Mr. Wilson invested money in two accounts. His total investment was $30,000. If one account pays 8% in interest and the other pays 4% in interest, how much does he have in each account if he earned a total of $1,960 in interest in 1 year?
He invested $
in the 8% account and $
in the 4% account.
I am completely lost with this process!!!

Answer by fractalier(6550)

Here's how we do this kind of problem...
Let us call x the amount he invested at 8%.
That means he invested 30000-x at 4%.
The set up looks like this:
.08x + .04(30000-x) = 1960
.08x + 1200 - .04x = 1960
.04x = 760
x = $19000 at 8% so that
30000-x = $11000 at 4%